Find three consecutive integers such that -7 times the sum of the first and third is 12 greater than the product of 10 and the opposite of the second.
Answer by mananth(16946)


Find three consecutive integers such that -7 times the sum of the first and third is 12 greater than the product of 10 and the opposite of the second.
let the integers be x,x+1,x+2
-7*(x+x+2) = 10*-(x+1)+12
-7(2x+2) = -10x-10+12
-14x-14 = -10x+2
4x=-16
x=-4
The integers are -4, -3,-2